# BODex: Scalable and Efficient Robotic Dexterous Grasp Synthesis Using Bilevel Optimization

### v2

@@ III Preliminaries
$\mathcal{F}_{i}$ =$=\left\{\mathbf{f}_{i}\in\mathbb{R}^{3}~{}|~{}0\leq f_{i,1}\leq 1,f_{i,2}^{2}+f_{i,3}^{2}\leq\mu^{2}f_{i,1}^{2}\right\}$ (1)
@@ IV-A Bilevel Optimization Formulation for Grasp Synthesis
By finding the optimal contact forces $\mathbf{y}_{j}$ for a desired wrench $\beta\mathbf{t}_{j}$, the grasp energy $Q_{j}(\mathbf{x})$ measures the difference between the desired wrench and the best-applied wrench $\Sigma_{i=1}^{m}\mathbf{G}_{i}\mathbf{f}_{j,i}$ at the grasp pose $\mathbf{x}$. To get a better grasp pose, the upper-level process performs gradient descent on $Q_{j}$, which is differentiable to $\mathbf{x}$ because $\mathbf{G}_{i}$ is differentiable to $\mathbf{x}$. The final grasp energy sums up several $Q_{j}$ with different $\mathbf{t}_{j}$ to encourage the applicable wrenches in multiple directions. For a force-closure grasp, $\{\mathbf{t}_{j}\}$ is set to be the six unit vectors along the positive and negative 3D force axes, with zero 3D torques, e.g., $[1,0,0,0,0,0]$ and $[-1,0,0,0,0,0]$. If a grasp can resist these six $\mathbf{t}_{j}$, it can resist the object gravity in any direction by a linear combination of these $\mathbf{t}_{j}$. Our formulation is also more flexible in customizing desired wrenches than the predefined primitive in TDG [19].
@@ V-A Simulation Environment and Object Assets
[…]
@@ V-B Evaluation Metrics
[…]
@@
[…]

### v3

@@ III Preliminaries
$\mathcal{F}_{i}$ $=\left\{\mathbf{f}_{i}\in\mathbb{R}^{3}~|~0\leq f_{i,1}\leq 1,f_{i,2}^{2}+f_{i,3}^{2}\leq\mu^{2}f_{i,1}^{2}\right\}$ (1)
@@ IV-A Bilevel Optimization Formulation for Grasp Synthesis
By finding the optimal contact forces $\mathbf{y}_{j}$ for a desired wrench $\beta\mathbf{t}_{j}$, the grasp energy $Q_{j}(\mathbf{x})$ measures the difference between the desired wrench and the best-applied wrench $\Sigma_{i=1}^{m}\mathbf{G}_{i}\mathbf{f}_{j,i}$ at the grasp pose $\mathbf{x}$. To get a better grasp pose, the upper-level process performs gradient descent on $Q_{j}$, which is differentiable to $\mathbf{x}$ because $\mathbf{G}_{i}$ is differentiable to $\mathbf{x}$. The final grasp energy sums up several $Q_{j}$ with different $\mathbf{t}_{j}$ to encourage the applicable wrenches in multiple directions. For a force-closure grasp, $\{\mathbf{t}_{j}\}$ is set to be the six unit vectors along the positive and negative 3D force axes, with zero 3D torques, e.g., $[1,0,0,0,0,0]$ and [1,0,0,0,0,0][-1,0,0,0,0,0]. If a grasp can resist these six $\mathbf{t}_{j}$, it can resist the object gravity in any direction by a linear combination of these $\mathbf{t}_{j}$. Our formulation is also more flexible in customizing desired wrenches than the predefined primitive in TDG [19].
@@ V-A Simulation Environment and Object Assets
Most parameters of the simulator and robot assets are retained as default, with only minimal adjustments: we set noslip_iterations to $2$ to prevent slow slippage, a characteristic of MuJoCo for solving inverse dynamics that is irrelevant to our task. The friction model uses a tangential coefficient of $0.6$ and a torsional coefficient of $0.02$, as recommended in [32]. Gravity uses $9.8\text{m/s}^{2}$. The object mass is set to $30\text{g}$.
@@ V-B Evaluation Metrics
Simulation Success Rate (SSR) (unit: $\%$) represents the percentage of successful grasps in simulation. To perform a grasp, the hand is initialized to the pre-grasp pose $\mathbf{x}_{p}$ and moves to the squeezed pose $\mathbf{x}_{s}$. Then, the object’s gravity is applied and we check whether the deviation in the object’s translation and rotation angle remains within $5\text{cm}$ and $15^{\circ}$, respectively, for more than $3$ seconds. Each grasp is tested across six orthogonal gravity directions and is regarded as a success only if it succeeds in all directions. Our environment evaluates $30.6$ grasps per second with $60$ threads on CPUs.
@@
First Variance Ratio (FVR) (unit: $\%$) is the ratio of the first eigenvalue in PCA, indicating the proportion of variance explained by the first principal component. Each grasp data point is a grasp pose $\mathbf{x}$ with the root rotation in 3D axis-angle format, normalized by setting the object to the identity pose.
@@ VII Conclusions
VIII Acknowledgment
This work was supported by Beijing Natural Science Foundation (Grant No.QY24042).